# RMT: Retentive Networks Meet Vision Transformers

### v1

@@ Results.
We compare RMT against many state-of-the-art models in Tab. 1. Results in the table demonstrate that RMT consistently outperforms previous models across all settings. Specifically, RMT-S achieves 84.1% Top1-accuracy with only 4.5 GFLOPs. RMT-B also surpasses iFormer [40] by 0.4% with similar FLOPs. Furthermore, our RMT-L model surpasses MaxViT-B [44] in top1-accuracy by 0.6% while using fewer FLOPs. Our RMT-T has also outperformed many lightweight models.
4.2 Object Detection and Instance Segmentation
Settings.
We adopt MMDetection [3] to implement RetinaNet [29] and Mask-RCNN [20]. We use the commonly used ”1×1$\times$1 ×” (12 training epochs) setting for the two strategies and ”$3\times+\mathrm{MS}$” for Mask-RCNN. Following [32], during training, images are resized to the shorter side of 800 pixels while the longer side is within 1333 pixels. We adopt the AdamW optimizer with a learning rate of 0.0001, weight decay of 0.05, and batch size of 16 to optimize the model. For ”1×1$\times$1 ×” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 8 and 11. While for ”$3\times+\mathrm{MS}$” schedule, the learning rate declines with the decay rate of 0.1 at the epoch 27 and 33.
Results.
Tab. 2 and Tab. 3 shows the results with RetinaNet and Mask R-CNN. The results demonstrate that our RMT performs best in all comparisons. For the RetinaNet framework, our RMT-T outperforms FAT-B2 by +1.1 AP, while S/B/L also perform better than other methods. As for the Mask R-CNN with ”1×1$\times$1 ×” schedule, RMT-L outperforms the recent InternImage-B by +1.8 box AP and +1.9 mask AP. For ”$3\times+\mathrm{MS}$” schedule, RMT-S outperforms InternImage-T for +1.6 box AP and +1.2 mask AP. All the above results tell that RMT outperforms its counterparts by evident margins.
@@ 4.3 Semantic Segmentation
Backbone Semantic FPN
Params(M) FLOPs(G) mIoU(%)
DAT-T [51] 32 198 42.6
CrossFormer-S [48] 34 221 46.0
UniFormer-S [27] 25 247 46.6
CSWin-T [9] 26 202 48.2
Shuted-S [39] 26 183 48.2
RMT-S 30 180 49.4
DAT-S 53 320 46.1
UniFormer-B [27] 54 350 47.7
CrossFormer-B [48] 56 331 47.7
CSWin-S [9] 39 271 49.2
RMT-B 57 294 50.4
DAT-B 92 481 47.0
CrossFormer-L 95 497 48.7
CSWin-B [9] 81 464 49.9
RMT-L 98 482 51.4
Table 4: Comparison with the state-of-the-art on ADE20K. The framework is SemanticFPN.
Backbone UperNet
Params(M) FLOPs(G) mIoU(%)
DAT-T [51] 60 957 45.5
NAT-T [18] 58 934 47.1
InternImage-T [49] 59 944 47.9
HorNet-T [38] 55 924 49.2
SMT-S [31] 50 935 49.2
RMT-S 56 937 49.8
DAT-S [51] 81 1079 48.3
SMT-B [31] 62 1004 49.6
HorNet-S [38] 85 1027 50.0
InterImage-S [49] 80 1017 50.2
CSWin-S [9] 65 1027 50.4
RMT-B 83 1051 51.5
Table 5: Comparison with the state-of-the-art on ADE20K. The framework is UperNet.
@@ Results.
The results of semantic segmentation can be found in Tab. 4 and Tab. 5. All the FLOPs are measured with the resolution of $512\times 2048$. All our models achieve the best performance in all comparisons. Specifically, our RMT-S exceeds Shunted-S for +1.2 mIoU with Semantic FPN. Moreover, our RMT-B outperforms the recent InternImage-S for +1.3 mIoU. All the above results demonstrate our model’s superiority in dense prediction.
@@ 4.4 Ablation Study
Model Params(M) FLOPs(G) Top1-acc(%)
RMT-T 14 2.5 82.4
w/o $\gamma$ 14 2.5 81.6
Table 6: Ablation study
$\gamma$ decay.
We verify the impact of explicit decay on the model, as shown in the Tab. 6. Explicit decay improves the performance of the model.
@@ 5 Conclusion
We propose RMT, a vision backbone that integrates retentive network and Vision Transformer. RMT introduces explicit decay related to distance, which brings spatial prior knowledge to visual models. The new mechanism is called Retentive Self-Attention (ReSA). To reduce the complexity of the model, RMT also employs a method that decomposes the ReSA into two axes. Extensive experiments validate the effectiveness of RMT, especially in downstream tasks such as object detection, where RMT demonstrates significant advantages. Our work is still in progress.

### v6

@@ Results.
Tab. 2, Tab. 3 and Tab. 4 show the results with different detection frameworks. The results demonstrate that our RMT performs best in all comparisons. For the RetinaNet framework, our RMT-T outperforms MPViT-XS by +1.3 AP, while S/B/L also perform better than other methods. As for the Mask R-CNN with “1×1$\times$1 ×” schedule, RMT-L outperforms the recent InternImage-B by +2.8 box AP and +1.9 mask AP. For “$3\times+\mathrm{MS}$” schedule, RMT-S outperforms InternImage-T for +1.6 box AP and +1.2 mask AP. Besides, regarding the Cascade Mask R-CNN, our RMT still performs much better than other backbones. All the above results tell that RMT outperforms its counterparts by evident margins.
@@ 4.3 Semantic Segmentation
Backbone Method Params(M) FLOPs(G) mIoU(%)
ResNet18 [21] FPN 15.5 32.2 32.9
PVTv2-B1 [54] FPN 17.8 34.2 42.5
VAN-B1 [17] FPN 18.1 34.9 42.9
EdgeViT-S [38] FPN 16.9 32.1 45.9
RMT-T FPN 17.0 33.7 46.4
DAT-T [58] FPN 32 198 42.6
RegionViT-S+ [3] FPN 35 236 45.3
CrossFormer-S [55] FPN 34 221 46.0
UniFormer-S [30] FPN 25 247 46.6
Shuted-S [44] FPN 26 183 48.2
RMT-S FPN 30 180 49.4
DAT-S [58] FPN 53 320 46.1
RegionViT-B+ [3] FPN 77 459 47.5
UniFormer-B [30] FPN 54 350 47.7
CrossFormer-B [55] FPN 56 331 47.7
CSWin-S [11] FPN 39 271 49.2
RMT-B FPN 57 294 50.4
DAT-B [58] FPN 92 481 47.0
CrossFormer-L [55] FPN 95 497 48.7
CSWin-B [11] FPN 81 464 49.9
RMT-L FPN 98 482 51.4
DAT-T [58] UperNet 60 957 45.5
NAT-T [19] UperNet 58 934 47.1
InternImage-T [56] UperNet 59 944 47.9
MPViT-S [29] UperNet 52 943 48.3
SMT-S [34] UperNet 50 935 49.2
RMT-S UperNet 56 937 49.8
DAT-S [58] UperNet 81 1079 48.3
SMT-B [34] UperNet 62 1004 49.6
HorNet-S [43] UperNet 85 1027 50.0
InterImage-S [56] UperNet 80 1017 50.2
MPViT-B [29] UperNet 105 1186 50.3
CSWin-S [11] UperNet 65 1027 50.4
RMT-B UperNet 83 1051 52.0
Swin-B [35] UperNet 121 1188 48.1
GC ViT-B [20] UperNet 125 1348 49.2
DAT-B [58] UperNet 121 1212 49.4
InternImage-B [56] UperNet 128 1185 50.8
CSWin-B [11] UperNet 109 1222 51.1
RMT-L UperNet 125 1241 52.8
Table 5: Comparison with the state-of-the-art on ADE20K.
@@
Model Params(M) FLOPs(G) Top1-acc(%) $AP^{b}$ $AP^{m}$ mIoU(%)
DeiT-S [49] 22 4.6 79.8
RMT-DeiT-S 22 4.6 81.7(+1.9)
Swin-T [35] 29 4.5 81.3 43.7 39.8 44.5
RMT-Swin-T 29 4.7 83.6(+2.3) 47.8(+4.1) 43.1(+3.3) 49.1(+4.6)
Swin-S [35] 50 8.8 83.0 45.7 41.1 47.6
RMT-Swin-S 50 9.1 84.5(+1.5) 49.5(+3.8) 44.2(+3.1) 51.0 (+3.4)
RMT-T 14.3 2.5 82.4 47.1 42.6 46.4
$\xrightarrow{}$ttention 14.3 2.5 81.6(-0.8) 44.6(-2.5) 40.7(-1.9) 43.9(-2.5)
$\xrightarrow{}$ate 15.6 2.7 Nan
w/o LCE 14.2 2.4 82.1 46.7 42.3 46.0
w/o CPE 14.3 2.5 82.2 47.0 42.4 46.4
w/o Stem 14.3 2.2 82.2 46.8 42.3 46.2
Table 6: Ablation study. We make a strict comparison among RMT, DeiT, and Swin-Transformer.
3rd stage FLOPs(G) Top1(%) FLOPs(G) mIoU(%)
MaSA-d 4.5 84.1 180 49.4
MaSA 4.8 84.1 246 49.7
Table 7: Comparison between decomposed MaSA (MaSA-d) and original MaSA.
[…]
Parallel 27 10.9 262
Chunklen_4 27 4.5 192
Chunklen_49 27 4.7 446 82.1
Recurrent 27 4.5 61
MaSA 27 4.5 876 84.1
Table 8: Comparison between MaSA and retention in RMT-S’s architecture.
[…]
BiFormer-T [75] 13 2.2 1602 81.4
CMT-XS [16] 15 1.5 1476 81.8
SMT-T [34] 12 2.4 636 82.2
RMT-T 14 2.5 1650 82.4
CMT-S [16] 25 4.0 848 83.5
MaxViT-T [51] 31 5.6 826 83.6
SMT-S [34] 20 4.8 356 83.7
BiFormer-S [75] 26 4.5 766 83.8
RMT-Swin-T 29 4.7 1192 83.6
RMT-S 27 4.5 876 84.1
SMT-B [34] 32 7.7 237 84.3
BiFormer-B [75] 57 9.8 498 84.3
CMT-B [16] 46 9.3 447 84.5
MaxViT-S [51] 69 11.7 546 84.5
RMT-Swin-S 50 9.1 722 84.5
RMT-B 54 9.7 457 85.0
SMT-L [34] 80 17.7 158 84.6
MaxViT-B [51] 120 23.4 306 84.9
RMT-L 95 18.2 326 85.5
Table 9: Comparison of inference speed among SOTA models.
@@ Results.
The results of semantic segmentation can be found in Tab. 5. All the FLOPs are measured with the resolution of $512\times 2048$, except the group of RMT-T, which are measured with the resolution of $512\times 512$. All our models achieve the best performance in all comparisons. Specifically, our RMT-S exceeds Shunted-S for +1.2 mIoU with Semantic FPN. Moreover, our RMT-B outperforms the recent InternImage-S for +1.8 mIoU. All the above results demonstrate our model’s superiority in dense prediction.
@@ 4.4 Ablation Study
Strict comparison with previous works.
In order to make a strict comparison with previous methods, we align RMT’s hyperparameters (such as whether to use hierarchical structure, the number of channels in the four stages of the hierarchical model, whether to use positional encoding and convolution stem, etc.) of the overall architecture with DeiT [49] and Swin [35], and only replace the Self-Attention/Window Self-Attention with our MaSA. The comparison results are shown in Tab. 6, where RMT significantly outperforms DeiT-S, Swin-T, and Swin-S.
MaSA.
We verify the impact of Manhattan Self-Attention on the model, as shown in the Tab. 6. MaSA improves the model’s performance in image classification and downstream tasks by a large margin. Specifically, the classification accuracy of MaSA is 0.8% higher than that of vanilla attention.
Softmax.
In RetNet, Softmax is replaced with a non-linear gating function to accommodate its various computational forms [46]. We replace the Softmax in MaSA with this gating function. However, the model utilizing the gating function cannot undergo stable training. It is worth noting that this does not mean the gating function is inferior to Softmax. The gating function may just not be compatible with our decomposed form or spatial decay.
LCE.
Local Context Enhancement also plays a role in the excellent performance of our model. LCE improves the classification accuracy of RMT by 0.3% and enhances the model’s performance in downstream tasks.
CPE.
Just like previous methods, CPE provides our model with flexible position encoding and more positional information, contributing to the improvement in the model’s performance in image classification and downstream tasks.
Convolutional Stem.
The initial convolutional stem of the model provides better local information, thereby further enhancing the model’s performance on various tasks.
Decomposed MaSA.
In RMT-S, we substitute the decomposed MaSA (MaSA-d) in the third stage with the original MaSA to validate the effectiveness of our decomposition method, as illustrated in Tab. 7. In terms of image classification, MaSA-d and MaSA achieve comparable accuracy. However, for semantic segmentation, employing MaSA-d significantly reduces computational burden while yielding similar result.
MaSA v.s. Retention.
As shown in Tab. 8, we replace MaSA with the original retention in the architecture of RMT-S. We partition the tokens into chunks using the method employed in Swin-Transformer [35] for chunk-wise retention. Due to the limitation of retention in modeling one-dimensional causal data, the performance of the vision backbone based on it falls behind RMT. Moreover, the chunk-wise and recurrent forms of retention disrupt the parallelism of the vision backbone, resulting in lower inference speed.
Inference Speed.
We compare the RMT’s inference speed with the recent best performing vision backbones in Tab. 9. Our RMT demonstrates the optimal trade-off between speed and accuracy.
@@ 5 Conclusion
In this work, we propose RMT, a vision backbone with explicit spatial prior. RMT extends the temporal decay used for causal modeling in NLP to the spatial level and introduces a spatial decay matrix based on the Manhattan distance. The matrix incorporates explicit spatial prior into the Self-Attention. Additionally, RMT utilizes a Self-Attention decomposition form that can sparsely model global information without disrupting the spatial decay matrix. The combination of spatial decay matrix and attention decomposition form enables RMT to possess explicit spatial prior and linear complexity. Extensive experiments in image classification, object detection, instance segmentation, and semantic segmentation validate the superiority of RMT.
Appendix A Architecture Details
Model Blocks Channels Heads Ratios Params(M) FLOPs(G)
RMT-T [2, 2, 8, 2] [64, 128, 256, 512] [4, 4, 8, 16] [3, 3, 3, 3] 14 2.5
RMT-S [3, 4, 18, 4] [64, 128, 256, 512] [4, 4, 8, 16] [4, 4, 3, 3] 27 4.5
RMT-B [4, 8, 25, 8] [80, 160, 320, 512] [5, 5, 10, 16] [4, 4, 3, 3] 54 9.7
RMT-L [4, 8, 25, 8] [112, 224, 448, 640] [7, 7, 14, 20] [4, 4, 3, 3] 95 18.2
RMT-DeiT-S [12] [384] [6] [4] 22 4.6
RMT-Swin-T [2, 2, 6, 2] [96, 192, 384, 768] [3, 6, 12, 24] [4, 4, 4, 4] 29 4.7
RMT-Swin-S [2, 2, 18, 2] [96, 192, 384, 768] [3, 6, 12, 24] [4, 4, 4, 4] 50 9.1
Table 10: Detailed Architectures of our models.
Our architectures are illustrated in the Tab. 10. For convolution stem, we apply five $3\times 3$ convolutions to embed the image into $56\times 56$ tokens. GELU and batch normalization are used after each convolution except the last one, which is only followed by batch normalization. $3\times 3$ convolutions with stride 2 are used between stages to reduce the feature map’s resolution. $3\times 3$ depth-wise convolutions are adopted in CPE. Moreover, $5\times 5$ depth-wise convolutions are adopted in LCE. RMT-DeiT-S, RMT-Swin-T, and RMT-Swin-S are models that we used in our ablation experiments. Their structures closely align with the structure of DeiT [49] and Swin-Transformer [35] without using techniques like convolution stem, CPE, and others.
Appendix B Experimental Settings
ImageNet Image Classification.
We adopt the same training strategy with DeiT [49] with the only supervision is the classification loss. In particular, our models are trained from scratch for 300 epochs. We use the AdamW optimizer with a cosine decay learning rate scheduler and 5 epochs of linear warm-up. The initial learning rate, weight decay, and batch size are set to 0.001, 0.05, and 1024, respectively. Our augmentation settings are RandAugment [8] (randm9-mstd0.5-inc1), Mixup [70] (prob=0.8), CutMix [69] (probe=1.0), Random Erasing [73] (prob=0.25) and Exponential Moving Average (EMA) [40]. The maximum rates of increasing stochastic depth [24] are set to 0.1/0.15/0.4/0.5 for RMT-T/S/B/L, respectively. For a more comprehensive comparison, we train two versions of the model. The first version uses only classification loss as the supervision, while the second version, in addition to the classification loss, incorporates token labeling introduced by [27] for additional supervision. Models using token labeling are marked with“*”.
COCO Object Detection and Instance Segmentation.
We apply RetinaNet [32], Mask-RCNN [22] and Cascaded Mask-CNN [2] as the detection frameworks to conduct experiments. We implement them based on the MMDetection [4]. All models are trained under two common settings:“1×1$\times$1 ×” (12 epochs for training) and“3×3$\times$3 ×+MS” (36 epochs with multi-scale augmentation for training). For the “1×1$\times$1 ×” setting, images are resized to the shorter side of 800 pixels. For the “3×3$\times$3 ×+MS”, we use the multi-scale training strategy and randomly resize the shorter side between 480 to 800 pixels. We apply AdamW optimizer with the initial learning rate of 1e-4. For RetinaNet, we use the weight decay of 1e-4 for RetinaNet while we set it to 5e-2 for Mask-RCNN and Cascaded Mask-RCNN. For all settings, we use the batch size of 16, which follows the previous works [35, 63, 64]
ADE20K Semantic Segmentation.
Based on MMSegmentation [7], we implement UperNet [59] and SemanticFPN [28] to validate our models. For UperNet, we follow the previous setting of Swin-Transformer [35] and train the model for 160k iterations with the input size of $512\times 512$. For SemanticFPN, we also use the input resolution of $512\times 512$ but train the models for 80k iterations.
Appendix C Efficiency Comparison
We compare the inference speed of RMT with other backbones, as shown in Tab. 11. Our models achieve the best trade-off between speed and accuracy among many competitors.
Model Params (M) FLOPs (G) Troughput (imgs/s) Top1 (%)
MPViT-XS [29] 11 2.9 1496 80.9
Swin-T [35] 29 4.5 1704 81.3
BiFormer-T [75] 13 2.2 1602 81.4
GC-ViT-XT [20] 20 2.6 1308 82.0
SMT-T [34] 12 2.4 636 82.2
RMT-T 14 2.5 1650 82.4
Focal-T [63] 29 4.9 582 82.2
CSWin-T [11] 22 4.3 1561 82.7
Eff-B4 [47] 19 4.2 627 82.9
MPViT-S [29] 23 4.7 986 83.0
Swin-S [35] 50 8.8 1006 83.0
SGFormer-S [15] 23 4.8 952 83.2
iFormer-S [45] 20 4.8 1051 83.4
CMT-S [16] 25 4.0 848 83.5
RMT-Swin-T 29 4.7 1192 83.6
CSwin-S [11] 35 6.9 972 83.6
MaxViT-T [51] 31 5.6 826 83.6
SMT-S [34] 20 4.8 356 83.7
BiFormer-S [75] 26 4.5 766 83.8
RMT-S 27 4.5 876 84.1
Model Params (M) FLOPs (G) Troughput (imgs/s) Top1 (%)
Focal-S [63] 51 9.1 351 83.5
Eff-B5 [47] 30 9.9 302 83.6
SGFormer-M [15] 39 7.5 598 84.1
SMT-B [34] 32 7.7 237 84.3
BiFormer-B [75] 57 9.8 498 84.3
RMT-Swin-S 50 9.1 722 84.5
MaxViT-S [51] 69 11.7 546 84.5
CMT-B [16] 46 9.3 447 84.5
iFormer-B [45] 48 9.4 688 84.6
RMT-B 54 9.7 457 85.0
Swin-B [35] 88 15.5 756 83.5
Eff-B6 [47] 43 19.0 172 84.0
Focal-B [63] 90 16.4 256 84.0
CSWin-B [11] 78 15.0 660 84.2
MPViT-B [29] 75 16.4 498 84.3
SMT-L [34] 80 17.7 158 84.6
SGFormer-B [15] 78 15.6 388 84.7
iFormer-L [45] 87 14.0 410 84.8
MaxViT-B [51] 120 23.4 306 84.9
RMT-L 95 18.2 326 85.5
Table 11: Comparison of inference speed.
Appendix D Details of Explicit Decay
We use different $\gamma$ for each head of the multi-head ReSA to control the receptive field of each head, enabling the ReSA to perceive multi-scale information. We keep all the $\gamma$ of ReSA’s heads within a certain range. Assuming the given receptive field control interval of a specific ReSA module is $[a,b]$, where both $a$ and $b$ are positive real numbers. And the total number of the ReSA module’s heads is $N$. The $\gamma$ for its $i$th head can be written as Eq. 8:
$\gamma_{i}=1-2^{-a-\frac{(b-a)i}{N}}$ (8)
For different stages of different backbones, we use different values of $a$ and $b$, with the details shown in Tab. 12.
Model $a$ $b$
RMT-T [2, 2, 2, 2] [6, 6, 8, 8]
RMT-S [2, 2, 2, 2] [6, 6, 8, 8]
RMT-B [2, 2, 2, 2] [7, 7, 8, 8]
RMT-L [2, 2, 2, 2] [8, 8, 8, 8]
RMT-DeiT-S [2] [8]
RMT-Swin-T [2, 2, 2, 2] [8, 8, 8, 8]
RMT-Swin-S [2, 2, 2, 2] [8, 8, 8, 8]
Table 12: Details about the $\gamma$ decay.